Let P QR be a triangle with coordinates P = (x1, y1, z1), Q = (x2, y2, z2) and R = (x3, y3, z3).

a) Prove that the triangle P QR lies on a plane. Find the equation of this plane.

b) Let (x, y, z) be a point in the plane containing the triangle. Prove that the det


x − x1 y − y1 z − z1

x − x2 y − y2 z − z2

x − x3 y − y3 z − z3

= 0.


1
Expert's answer
2021-09-24T12:12:18-0400

a) Find the equation of a plane passing through three non-collinear points "(x_1 , y_1 , z _1), (x_2 , y_2 , z_ 2)" and "(x_3 , y_3 , z _3)."



"ax_1+by_1+cz_1=d""ax_2+by_2+cz_2=d""ax_3+by_3+cz_3=d"

The system has unique solution if and only if



"\begin{vmatrix}
 x_1 & y_1 & z_1 \\
 x_2 & y_2 & z_2 \\
 x_3 & y_3 & z_3 \\
\end{vmatrix}\not=0"

(b)

Let "(x, y, z)" be a point in the plane containing the triangle. Then



"ax+by+cz=d""ax_1+by_1+cz_1=d""ax_2+by_2+cz_2=d""ax_3+by_3+cz_3=d"

Then



"a(x-x_1)+b(y-y_1)+c(z-z_1)=0""a(x_2-x_1)+b(y_2-y_1)+c(z_2-z_1)=0""a(x_3-x_1)+b(y_3-y_1)+c(z_3-z_1)=0"


System has nonzero solutions for "(a,b,c)" if



"D=\begin{vmatrix}
 x-x_1 & y-y_1 & z-z_1 \\
 x_2-x_1 & y_2-y_1 & z_2-z_1 \\
 x_3-x_1 & y_3-y_1 & z_3-z_1 \\
\end{vmatrix}=0"